Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (A. Kraus at Thu Feb 15 12:06:20 MET 2001) Success

Noto (C.Trigilio at Thu Feb 15 18:52:55 MET 2001) Minor failures
Lost data from 00:04 to 00:23 for an anomalous stop of the tape drive.

Westerbork (Tony Foley at Fri Feb 16 16:36:00 MET 2001) Minor failures
A section of this was outside our hour angle range and so was replaced by local calibration.

Jodrell_Bank (Alastair Gunn at Wed Feb 21 16:09:10 MET 2001) Minor failures
Small loss of data due to FS crash between 02:26 and 03:14 UT. Most data should be good even during this period. No other known problems.